OpenStudy (anonymous):

Sinusoidal means looking like a basic sine graph. By how many degrees do you have to shift sin x = y to the right to make it look like the graph of cos x = y?

OpenStudy (anonymous):

\[ \sin(x) = \cos(\pi/2-x) \]

OpenStudy (anonymous):

Usually, we have the equation: \[sin\left(x+\frac{\pi}{2}\right)=sin\left(x-\frac{3\pi}{2}\right)=cos(x)\] So you either need to shift \(sin(x)\): \(\frac{\pi}{2}\) To the left or: \(\frac{3\pi}{2}\) to the right

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Anytime